OpenStudy (anonymous):

Trigonometry Question -- Need urgent help ! Thank You <3

OpenStudy (anonymous):

\[\frac {\sqrt{3} \cos 33 - \sin 23 }{2} = \cos 53\]

OpenStudy (mathteacher1729):

what are you supposed to do with this?

OpenStudy (anonymous):

We need to prove the same!

OpenStudy (anonymous):

thats \(\cos 23\) i think ha?

OpenStudy (anonymous):

Nopes! It's cos 33 itself! :(

OpenStudy (mathteacher1729):

The expression you have written is not true. The left side of the equal sign does not equal the right.

OpenStudy (mathteacher1729):

(sqrt(3)*cos(23)-sin(23))/2 = cos(53) is true.

OpenStudy (anonymous):

Oh! Thanx! Maybe i noted the question wrong! :/ :(

OpenStudy (anonymous):

cos 33 = 0.8387 sin 23 = 0.3907 cos 53 = 0.6018 sqrt(3) = 1.732 >>> (1.732 * 0.8387 - 0.3907)/2.0 0.5309642

OpenStudy (anonymous):